a man travels from a to b at 4km/h and from b to a at 6km/h.the total journey takes 45 minutes .find the distance travelled?

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

a man travels from a to b at 4km/h and from b to a at 6km/h.the total journey takes 45 minutes .find the distance travelled?
---------
avg speed for the round trip is 2*4*6/(4+6) = 4.8 km/hr
RT distance = 4.8*0.75 = 3.6 km

Answer by rothauserc(4718)

This is a rate * time = distance problem
Let t1 be time to go from a to b and t2 be time to go from b to a, then
4*t1 = 6*t2, also
Note that 45 minutes is 3/4 hour or .75 hour, then
t1 + t2 = .75
Solve for t1
t1 = .75 - t2
Substitute for t1 in first equation
4*(.75 - t2) = 6*t2
3 -4t2 = 6t2
10t2 = 3
Note that 3 hours is 180 minutes
10t2 = 180
t2 = 18
t1 = 45 -18 = 27 minutes
Now distance from a to b is 4*27/60 = 27/15 = 9/5 = 1.8 km
Total distance is 2 * 1.8 = 3.6 km